40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15181     Entry time: Fri Jan 31 16:04:30 2020
Author: gautam 
Type: Update 
Category: IOO 
Subject: Input pointing drift 

One factor which hampers locking efforts is the apparent drift of the input beam into the IFO. Over timescales of ~1 hour, I have noticed that the spot on the AS camera drifts significantly (~1 spot size) in pitch. The IPPOS QPD bears out this observation, see Attachment #1. The IMC WFS control signals do not show a correlated drift, hence my claim that the TTs are to blame.

I am able to correct this misalignment by moving TT1 in pitch (see Attachment #2, which shows some signals from a ~1 hour PRMI lock, during which time the pointing drifted, and I corrected it by moving TT1 pitch). Assuming the problem is purely TT1 pitch drifting, this corresponds to 3mm / 6m ~500urad of shift in 1 hour - seems very large. The fact that the drift is only present in pitch and doesn't really show up in yaw makes me think the problem is likely mechanical (unless the voltage to the top two coils is drifting relative to the bottom, but no LR drift, which would be very coincidental). At the moment, this is just an annoyance, but it'd be good for this problem to be fixed.

I haven't yet figured out how to make ndscope export these plots to SVG preserving the dark color theme, hence the weird light axes...

Attachment 1: IPdrift.pdf  168 kB  | Hide | Hide all
IPdrift.pdf
Attachment 2: IPdrift_PRMI.pdf  291 kB  | Hide | Hide all
IPdrift_PRMI.pdf
ELOG V3.1.3-